# Dual-Stream Diffusion for World-Model Augmented Vision-Language-Action Model

### v1

@@ Abstract
Recently, augmenting Vision-Language-Action models (VLAs) with world modeling has shown promise in improving robotic policy learning. However, it remains challenging to jointly predict next-state observations and action sequences because of the inherent difference between the two modalities. To address this, we propose DUal-STream diffusion (DUST), a world-model augmented VLA framework that handles the modality conflict and enhances the performance of VLAs across diverse tasks. Specifically, we propose a multimodal diffusion transformer architecture that explicitly maintains separate modality streams while still enabling cross-modal knowledge sharing. In addition, we introduce independent noise perturbations for each modality and a decoupled flow-matching loss. This design enables the model to learn the joint distribution in a bidirectional manner while avoiding the need for a unified latent space. Based on the decoupling of modalities during training, we also introduce a joint sampling method that supports test-time scaling, where action and vision tokens evolve asynchronously at different rates. Through experiments on simulated benchmarks such as RoboCasa and GR-1, DUST achieves up to 6% gains over baseline methods, while our test-time scaling approach provides an additional 2–5% boost. On real-world tasks with the Franka Research 3, DUST improves success rates by 13%, confirming its effectiveness beyond simulation. Furthermore, pre-training on action-free videos from BridgeV2 yields significant transfer gains on RoboCasa, underscoring DUST’s potential for large-scale VLA pretraining.
@@ 1 Introduction
Vision-Language-Action models (VLAs) have recently emerged as promising approaches for learning general-purpose robotic policies (Black et al., 2025; NVIDIA et al., 2025). These models extend Vision-Language Models (VLMs), which are pretrained on internet-scale multimodal datasets, by further finetuning them on robotics datasets. This enables VLAs to generate actions that generalize to novel objects, scenes, and instructions (Zawalski et al., 2024). Despite strong perception grounding and instruction following, they fail to model how actions affect the environment, and thus lack an explicit understanding of underlying physical processes. To address this, recent works have augmented VLAs with world modeling objectives, which train models to predict future visual observations alongside actions (Guo et al., 2024; Zheng et al., 2025; Liang et al., 2025). This joint prediction of the two modalities enables the model to effectively capture the latent dynamics that govern both actions and their visual results, resulting in improved performance and generalization.
@@
Existing works that augment VLAs with world modeling such as PAD (Guo et al., 2024) and EnerVerse (Huang et al., 2025) utilize unified joint diffusion model structures (See Figure 1(a)), where the two modalities are concatenated together and modeled with a single unified model. However, this design implicitly assumes a shared latent space between modalities. As a result, the model must jointly optimize two fundamentally different objectives, where action trajectory predictions require low-dimensional, temporally smooth outputs, while future visual observations require high-dimensional, spatially structured outputs. In contrast, approaches such as Video Policy (Liang et al., 2025) and Video Prediction Policy (Hu et al., 2025) adopt a causal diffusion design (See Figure 1(b)) that separates the two modalities into distinct models connected through unidirectional conditioning. While this respects modality-specific structure, it limits information flow to a single direction, preventing bidirectional knowledge transfer. Together, these two design choices highlight a trade-off between cross-modal integration and modality-specific fidelity. To bridge these contrasting approaches, we propose DUal-STream Diffusion (DUST), a VLA architecture that preserves distinct modality streams while facilitating information exchange across them (See Figure 1(c)).
DUST introduces a dual-stream, multimodal diffusion transformer that maintains separate pathways for action and vision tokens, yet enables interaction through shared cross-modal attention layers. On top of this architecture, we introduce a decoupled diffusion training algorithm that applies independent noising to each modality and optimizes a modality-specific flow-matching loss. This formulation allows actions and observations to be learned according to their distinct statistical structures, while enabling the model to capture cross-modal causal dependencies. Building on the decoupled noising process, we further propose a joint sampling strategy that enables test-time scaling through asynchronous denoising of the two modalities.
@@
We evaluate DUST across simulated, real-world, and transfer learning scenarios to assess its effectiveness and scalability. In simulation, DUST surpasses baselines such as GR00T-N1.5 (NVIDIA et al., 2025) and FLARE (Zheng et al., 2025) on the RoboCasa and GR-1 benchmarks, with success rate improvements of up to 5% and 6%, respectively. On a real Franka Research 3 arm, it achieves the highest success rates across diverse pick-and-place tasks, outperforming baselines by more than 12%. When pretrained on action-free video from BridgeV2, DUST exhibits significant gains during downstream fine-tuning on RoboCasa, which demonstrates that it can leverage large-scale passive video for data-efficient policy learning. The asynchronous joint diffusion sampling strategy also proves effective at test time, providing an additional 2–6% boost over naive sampling approaches.
@@
Figure 1: Architectures for world modeling augmentations to VLAs. (a) Unified Joint Diffusion concatenates action and vision tokens and generates both with a single model. (b) Causal Diffusion uses separate models with one-way conditioning. (c) Dual-Stream Diffusion maintains separate streams for each modality while enabling cross-modal knowledge transfer through shared attention.
@@ 2 Related Works
Vision-Language-Action models (VLAs).  Vision–Language–Action models (VLAs) have recently emerged as a promising paradigm for general-purpose robot policy learning, extending Vision–Language Models (VLMs) pretrained on internet-scale multimodal datasets. Building on the strong representational capacity of VLMs (Dai et al., 2023; Team, 2024; Xiao et al., 2024), VLA architectures adapt them for robotics by either generating actions autoregressively (Kim et al., 2024; Brohan et al., 2023; Wu et al., 2024; Cheang et al., 2024) or employing diffusion-based action experts (Black et al., 2025; NVIDIA et al., 2025). In this work, we adopt the diffusion modeling formulation for action generation. Beyond these designs, recent extensions explore cross-embodiment latent action spaces (Ye et al., 2025; Bu et al., 2025) and reasoning-driven architectures for complex task execution (Zawalski et al., 2024). Despite these advances, most approaches emphasize imitation-based action distribution learning without explicitly modeling how actions influence future states. In contrast, our framework integrates a world-modeling objective that captures physical dynamics, enabling more grounded and effective action generation.
World-modeling for robotic policy learning.  Prior work has augmented VLAs with world-modeling objectives that generate future states alongside action generation. One line of research, exemplified by PAD (Guo et al., 2024) and EnerVerse (Huang et al., 2025), employs unified architectures that jointly model future images and actions through diffusion (Figure 1(a)). UWM (Zhu et al., 2025) extends this approach with modality-specific time schedules, while FLARE (Zheng et al., 2025) introduces implicit world modeling by aligning mid-level features to future image embeddings instead of directly diffusing them. UVA (Li et al., 2025a) embeds both modalities into a shared latent space, followed by modality-specific decoders that reconstruct their native representations. A complementary line of work, including Video Policy (Liang et al., 2025) and Video Prediction Policy (Hu et al., 2025), adopts disjoint architectures that allow only unidirectional conditioning between modalities (Figure 1(b)).
Another key design choice concerns how future states are represented. A common approach used in PAD (Guo et al., 2024), PIDM (Tian et al., 2025), and This&That (Wang et al., 2024) is to reconstruct the next RGB observation directly after executing the generated action segment. In contrast, methods such as DINO-WM (Zhou et al., 2024) and FLARE (Zheng et al., 2025) replace raw image prediction with the generation of future observation embeddings derived from pretrained encoders like DINO-V2 (Oquab et al., 2023) and Q-Former (Li et al., 2023). We adopt this embedding-based strategy, as it emphasizes the semantic structure of future states while avoiding the need to reproduce pixel-level details, which is information that is often irrelevant for downstream control, yet costly to model.
@@ 3 Preliminaries
Problem setup.  Let $\mathcal{D}=\{T_{1},T_{2},$...\} be the dataset composed of expert demonstration trajectories, where each trajectory $T_{i}=\{I,\,\{(O_{t},A_{t})\}_{t=0}^{L}\}$ consists of task instruction $I$ and observations $O_{t}$ and action sequences $A_{t}$. Specifically, we denote the observations at timestep $t$ as $O_{t}=(o_{t}^{v},o_{t}^{s})$, where $o_{t}^{v}$ is the visual observation and $o_{t}^{s}$ is the robot proprioceptive state. Actions are grouped in chunks (Zhao et al., 2023; Chi et al., 2023) such that $A_{t}=(a_{t},a_{t+1},...,a_{t+k-1})$ where $k$ is the size of chunk. Our goal is to train a model that predicts $A_{t}$ given the observations $O_{t}$ and instruction $I$.
Vision-Language-Action model (VLA). In developing a VLA model to solve this problem, we follow common practice introduced in recent diffusion-based VLA models (Black et al., 2025; NVIDIA et al., 2025). Specifically, we use a pretrained Vision-Language model (VLM; Li et al. 2025b) to extract high-level semantic information from the image observations and text instruction. Then, the extracted representations are used as conditions for the action expert, e.g., through cross-attention layers in diffusion transformers (DiT; Peebles & Xie 2022) for action prediction.
@@
$\mathcal{L}_{\textrm{FM}}(\theta)=\mathbb{E}_{A_{t}^{\tau},\tau}\Big[\big\|V_{\theta}(A_{t}^{\tau},\Phi_{t},o_{t}^{s})-(A_{t}-\epsilon)\big\|^{2}\Big]\text{,}$ (1)
@@
World-modeling.  The objective of world modeling is to learn predictive representations of future states. We consider predicting future image observation $o_{t+k}^{v}$, which is obtained by executing the action chunk $A_{t}$ of length $k$. However, direct pixel-level prediction may lead to emphasis on learning of high-frequency visual details that are irrelevant to low-level control and hinder the learning of meaningful physical dynamics. To this end, we instead predict the representation of the future visual observation, which we obtain by re-using the vision encoder in our VLM to embed the future image. We denote $\tilde{o}_{t+k}$ to be the future image embedding, and our world modeling goal is to predict this embedding conditioned on VLM features $\Phi_{t}$ and proprioceptive state $o_{t}^{s}$.
@@ 4 Method
Figure 2: Dual-Stream diffusion (DUST) architecture. Our architecture has (1) VLM model $\text{VLM}\phi(\cdot)$ that processes current observation and task instruction to produce semantic representations, and (2) diffusion model $\pi_{\theta}$ which conditions on these representations to generate actions and future observation embeddings.
In this section, we present the Dual-STream diffusion (DUST) model, our framework designed for joint world modeling and action prediction. The core challenge we address is the inherent conflict within joint modeling of the two modalities, actions and future observations, which have fundamentally different statistical properties. Our method systematically resolves this conflict through three key contributions. We first introduce the DUST architecture (Section 4.1), which utilizes a multi-modal diffusion transformer to maintain modality-specific pathways while enabling cross-modal information exchange. We then detail our decoupled training algorithm (Section 4.2), which employs independent noise levels for each modality during training to optimize a joint objective. Finally, we describe a novel joint sampling strategy (Section 4.3) that supports test-time scaling by evolving the two modalities at different rates.
@@ 4.1 DUST architecture
This input is processed by a stack of Multi-Modal Diffusion Transformer (MMDiT) blocks. Critically, within each MMDiT block, the action and vision token streams are propagated through separate pathways. They are concatenated only temporarily during the shared cross-modal attention layer, which facilitates information exchange, and are immediately split back into their respective streams for all other operations. To further decouple their dynamics and directly support our training objective (described in Section 4.2), each stream receives its own distinct timestep embedding via Adaptive LayerNorm (AdaLN;Peebles & Xie (2022)). After traversing the shared MMDiT layers, the two streams are routed into their own modality-specific DiT blocks for several layers of fine-grained, specialized denoising. This final stage allows the vision pathway to focus on reconstructing a semantically consistent future embedding, while the action pathway refines the low-level motor control trajectories, thereby improving the joint modeling of both control and world dynamics.
@@ 4.2 Joint training algorithm
By sampling separate timesteps, we allow for causal dependencies to be learned between the modalities. For example, the model might be asked to predict a nearly-clean future observation ($\tau_{o}\approx 1$) from a completely noisy action ($\tau_{A}\approx 0$). This forces it to learn the inverse relationship, effectively answering: "What action must have been taken to achieve this future state?" Conversely, the model might be given a nearly-clean action ($\tau_{A}\approx 1$) and be required to denoise a very noisy future observation ($\tau_{o}\approx 0$). This trains the forward causal relationship: "What will the future state look like as a result of this action?" This independence allows each modality to be noised and denoised according to its own properties, and this varied training across all combinations of noise levels is precisely what enables the model to capture the deep causal relationships between them, all managed through the shared-attention architecture.
Decoupled noise scheduling.  Let the two modalities be the action chunk $A_{t}\in\mathbb{R}^{k\times d_{A}}$ and the future observation embedding $\tilde{o}_{t+k}\in\mathbb{R}^{d_{o}}$, where $d_{A}$ and $d_{o}$ are the dimensions of the action space and future image embedding, respectively. During training, we sample timesteps independently, with $\tau_{A}\in[0,1]$ for actions and $\tau_{o}\in[0,1]$ for future visual observations. Let $\epsilon_{A},\epsilon_{o}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ be sampled Gaussian noise, with which we noise $A_{t}$ and $\tilde{o}_{t+k}$, giving the noisy action sequences and noisy future observations as $A_{t}^{\tau_{A}}=\tau_{A}A_{t}+(1-\tau_{A})\epsilon_{A}$ and $\tilde{o}_{t+k}^{\tau_{o}}=\tau_{o}\tilde{o}_{t+k}+(1-\tau_{o})\epsilon_{o}$, respectively. The diffusion model $V_{\theta}$ predicts the velocity field of each modality, conditioned on the VLM feature $\Phi_{t}$. Let us denote $V_{\theta}(A_{t}^{\tau_{A}},\tilde{o}_{t+k}^{\tau_{o}},\Phi_{t},o_{t}^{s})=[V_{\theta}^{A},V_{\theta}^{o}]$ be the outputs of diffusion model. Then, the training objective for each action and image observations (i.e., world-modeling) are given as follows:
@@
where $\lambda_{\textrm{WM}}>0$ is a weighting hyperparameter for world modeling loss.
Figure 3: Overview of vision-action joint sampling. During inference, we sample over $N_{A}$ steps for action tokens and $N_{o}=q\times N_{A}$ steps for vision tokens. The global timestep advances by $\Delta\tau_{o}=1/N_{o}$, and action tokens are updated only every $q$ steps in $\Delta\tau_{A}=1/N_{A}$ strides. Default $q$ value is 1, and increasing it allows test-time scaling.

### v2

@@ Abstract
Recently, augmenting vision-language-action models (VLAs) with world-models has shown promise in robotic policy learning. However, it remains challenging to jointly predict next-state observations and action sequences because of the inherent difference between the two modalities. To address this, we propose dual-stream diffusion (DUST), a world-model augmented VLA framework that handles the modality conflict and enhances the performance of VLAs across diverse tasks. Specifically, we propose a multimodal diffusion transformer architecture that explicitly maintains separate modality streams while enabling cross-modal knowledge sharing. In addition, we propose training techniques such as independent noise perturbations for each modality and a decoupled flow matching loss, which enables the model to learn the joint distribution in a bidirectional manner while avoiding the need for a unified latent space. Furthermore, based on the decoupled training framework, we introduce a sampling method where we sample action and vision tokens asynchronously at different rates, which shows improvement through inference-time scaling. Through experiments on simulated benchmarks such as RoboCasa and GR-1, DUST achieves up to 6% gains over a standard VLA baseline and implicit world-modeling methods, with our inference-time scaling approach providing an additional 2-5% gain on success rate. On real-world tasks with the Franka Research 3, DUST outperforms baselines in success rate by 13%, confirming its effectiveness beyond simulation. Lastly, we demonstrate the effectiveness of DUST in large-scale pretraining with action-free videos from BridgeV2, where DUST leads to significant gain when transferred to the RoboCasa benchmark.
@@ 1 Introduction
Vision-language-action models (VLAs) have recently emerged as promising approaches for learning general-purpose robotic policies (Black et al., 2025; NVIDIA et al., 2025; Brohan et al., 2023; Li et al., 2023b; Kim et al., 2024; Luo et al., 2025; Shukor et al., 2025). Specifically, VLAs are built upon vision-language models (VLMs), which are pretrained on internet-scale multimodal datasets and excel in visual and textual understanding. Then, VLAs leverage VLM features to generate actions, through action experts (e.g., diffusion policy (Chi et al., 2023)), that generate robotic actions given the current observation and text instruction. As such, VLAs are capable of generating precise actions that generalize to novel objects, scenes, and instructions (Zawalski et al., 2024). However, despite strong perceptual grounding and instruction following capabilities, VLAs often fail to model how actions affect the environment, and fall short in terms of explicit understanding of underlying physical processes (Guo et al., 2024).
@@
To address this, recent works have augmented VLAs with world-modeling objectives, which train models to predict future visual observations together with actions (Guo et al., 2024; Zheng et al., 2025; Liang et al., 2025). Through learning the joint distribution of the two modalities, it enables the models to effectively capture the latent dynamics that govern both actions and their visual results, improving performance and generalization. Previous works utilized unified joint diffusion model structures (e.g., see Figure 1(a)), where the two modalities are concatenated together and modeled with a single unified model (Guo et al., 2024; Huang et al., 2025). However, their design relies on the implicit assumption of the existence of a shared latent space between the modalities. As a result, the model often suffers from mismatch between modalities, where action predictions require low-dimensional, temporally smooth outputs, while future visual observations require high-dimensional, spatially structured outputs. Other approaches adopt a causal design (e.g., see Figure 1(b)), which separates the modalities into distinct models with uni-directional conditioning (Liang et al., 2025; Hu et al., 2025). While this design can handle modality-specific structure, the design inherently limits information flow to a single direction, and prevents bidirectional knowledge transfer. As such, designing world-models and action prediction models remains a challenge due to the trade-off between cross-modal integration and modality-specific fidelity.
Contribution. To bridge these contrasting approaches, we propose dual-stream diffusion (DUST), a VLA architecture that preserves distinct modality streams while facilitating information exchange across them (see Figure 1(c)). DUST employs a multimodal diffusion transformer (MMDiT) (Esser et al., 2024) that maintains separate token streams for actions and future visual observations, each with its own timestep embedding and normalization. The two modality streams interact through shared cross-modal attention layers, allowing bidirectional information flow without collapsing modalities into a single latent space. On top of this architecture, we introduce a decoupled diffusion training algorithm that applies independent noising schedules to each modality, enabling the model to learn causal relationships between them under various noise configurations (Chen et al., 2025; Rojas et al., 2025). The network is optimized via modality-specific flow matching losses, allowing actions and observations to evolve according to their respective statistical structures. Finally, we introduce a unique sampling strategy for DUST that jointly samples action and visual observations. Specifically, in order to handle the difference between the modalities, we introduce asynchronous denoising, where we take diffusion steps on the high-dimensional vision tokens more frequently than the low-dimensional action tokens. As a result, our approach yields scalable test-time scaling that effectively balances efficiency and accuracy.
@@
We evaluate the effectiveness and scalability of DUST through extensive evaluations on simulated, real-world, and transfer learning scenarios. To analyze DUST’s pretraining performance, we freeze the backbone VLM (Li et al., 2025b) and train the diffusion-based action expert (Chi et al., 2023) across all experiments. In simulation benchmarks, DUST outperforms baselines such as standard VLAs (e.g., GR00T-N1.5 (NVIDIA et al., 2025)) and implicit world-modeling approaches (e.g., FLARE (Zheng et al., 2025)) on the RoboCasa and GR-1 benchmarks, achieving success rate gains of 5% and 6%, respectively. When evaluating on a real Franka Research 3 arm, DUST achieves the highest success rates across diverse pick-and-place tasks, outperforming baselines by over 12%, and demonstrates robust real-world performance and physically consistent predictions across environments. Lastly, we leverage DUST in pretraining with action-free videos (e.g., BridgeV2 (Walke et al., 2023)), and show that DUST exhibits significant gains when transferred to downstream tasks, such as the RoboCasa benchmark. The asynchronous joint diffusion sampling strategy also proves effective at test-time, providing an additional 2–6% boost over naive sampling approaches.
@@
Figure 1: Architectures of world-model augmented VLAs. (a) Unified Joint Diffusion concatenates action and vision tokens and generates both with a single model. (b) Causal Diffusion uses separate models with one-way conditioning. (c) Dual-Stream Diffusion (ours) maintains separate streams for each modality while enabling cross-modal knowledge transfer through shared attention.
@@ 2 Related Works
Vision-language-action models (VLAs).  VLAs have recently emerged as a promising paradigm for general-purpose robot policy learning, extending vision–language models (VLMs) pretrained on internet-scale multimodal datasets. Building on the strong representational capacity of VLMs (Dai et al., 2023; Team, 2024; Xiao et al., 2024), VLA architectures adapt them for robotics by either generating actions autoregressively (Kim et al., 2024; Brohan et al., 2023; Wu et al., 2024; Cheang et al., 2024) or employing diffusion-based action experts (Black et al., 2025; NVIDIA et al., 2025). In this work, we adopt the diffusion modeling formulation for action generation. Beyond these designs, recent extensions explore cross-embodiment latent action spaces (Ye et al., 2025; Bu et al., 2025) and reasoning-driven architectures for complex task execution (Zawalski et al., 2024). Despite these advances, most approaches emphasize imitation-based action distribution learning without explicitly modeling how actions influence future states. In contrast, our framework integrates a world-modeling objective that captures physical dynamics, enabling more grounded and effective action generation.
World-modeling for robotic policy learning.  Prior work has augmented VLAs with world-modeling objectives that generate future states alongside action generation. One line of research, exemplified by PAD (Guo et al., 2024) and EnerVerse (Huang et al., 2025), employs unified architectures that jointly model future images and actions through diffusion (Figure 1(a)). UWM (Zhu et al., 2025) extends this approach with modality-specific time schedules, while FLARE (Zheng et al., 2025) introduces implicit world-modeling by aligning mid-level features to future image embeddings instead of directly diffusing them. UVA (Li et al., 2025a) embeds both modalities into a shared latent space, followed by modality-specific decoders that reconstruct their native representations. A complementary line of work, including Video Policy (Liang et al., 2025) and Video Prediction Policy (Hu et al., 2025), adopts disjoint architectures that allow only unidirectional conditioning between modalities (Figure 1(b)).
Another key design choice concerns how future states are represented. A common approach used in PAD (Guo et al., 2024), PIDM (Tian et al., 2025), and This&That (Wang et al., 2024) is to reconstruct the next RGB observation directly after executing the generated action segment. In contrast, methods such as DINO-WM (Zhou et al., 2024) and FLARE (Zheng et al., 2025) replace raw image prediction with the generation of future observation embeddings derived from pretrained encoders like DINO-V2 (Oquab et al., 2023) and Q-Former (Li et al., 2023a). We adopt this embedding-based strategy, as it emphasizes the semantic structure of future states while avoiding the need to reproduce pixel-level details, which is information that is often irrelevant for downstream control, yet costly to model.
@@ 3 Preliminaries
Problem setup.  Let $\mathcal{D}=\{T_{1},T_{2},$...\} be the dataset composed of expert demonstration trajectories, where each trajectory $T_{i}=\{I,\,\{(O_{t},A_{t})\}_{t=0}^{L}\}$ consists of task instruction $I$ and observations $O_{t}$ and action sequences $A_{t}$. Specifically, we denote the observations at timestep $t$ as $O_{t}=(o_{t}^{v},o_{t}^{s})$, where $o_{t}^{v}$ is the visual observation and $o_{t}^{s}$ is the robot proprioceptive state. Actions are grouped in chunks (Zhao et al., 2023; Chi et al., 2023) such that $A_{t}=(a_{t},a_{t+1},...,a_{t+k-1})$ where $k$ is the chunk length. Our goal is to train a model that predicts $A_{t}$ given observations $O_{t}$ and instruction $I$.
Vision-language-action model (VLA). In developing a VLA model to solve this problem, we follow common practice introduced in recent diffusion-based VLA models (Black et al., 2025; NVIDIA et al., 2025). Specifically, we use a pretrained vision-language model (VLM; Li et al. 2025b) to extract high-level semantic information from the image observations and text instruction. Then, the extracted representations are used as conditions for the action expert through cross-attention layers in a diffusion transformer (DiT; Peebles & Xie 2022) during action prediction.
@@
$\mathcal{L}_{\textrm{FM}}(\theta)=\mathbb{E}_{A_{t}^{\tau},\tau}\Big[\big\|V_{\theta}(\Phi_{t},A_{t}^{\tau},o_{t}^{s})-(A_{t}-\epsilon)\big\|^{2}\Big]\text{,}$ (1)
@@
World-modeling.  The objective of world-modeling is to learn predictive representations of future states. We consider predicting future image observation $o_{t+k}^{v}$, which is obtained by executing the action chunk $A_{t}$ of length $k$. However, direct pixel-level prediction may lead to emphasis on learning of high-frequency visual details that are irrelevant to low-level control and hinder the learning of meaningful physical dynamics. To this end, we instead predict the representation of the future visual observation, which we obtain by re-using the vision encoder in our VLM to embed the future image. We denote $\tilde{o}_{t+k}$ to be the future image embedding, and our world-modeling goal is to predict this embedding conditioned on VLM features $\Phi_{t}$ and proprioceptive state $o_{t}^{s}$.
@@ 4 Method
Figure 2: Dual-stream diffusion (DUST) architecture. Our architecture has a (1) VLM model $\text{VLM}\phi(\cdot)$ that processes current observation and task instruction to produce semantic representations, and a (2) diffusion model $\pi_{\theta}$ which conditions on these representations to generate actions and future observation embeddings.
In this section, we present the dual-stream diffusion (DUST) model, our framework designed for joint world-modeling and action prediction. The core challenge we address is the inherent conflict within joint modeling of the two modalities, actions and future observations, which have fundamentally different statistical properties. Our method systematically resolves this conflict through three key contributions. We first introduce the DUST architecture (Section 4.1), which utilizes a multimodal diffusion transformer to maintain modality-specific pathways while enabling cross-modal information exchange. We then detail our decoupled training algorithm (Section 4.2), which employs independent noise levels for each modality during training to optimize a joint objective. Finally, we describe a novel joint sampling strategy (Section 4.3) that supports test-time scaling by evolving the two modalities at different rates.
@@ 4.1 DUST architecture
This input is processed by a stack of multimodal diffusion transformer (MMDiT) blocks. Critically, within each MMDiT block, the action and vision token streams are propagated through separate pathways. They are concatenated only temporarily during the shared cross-modal attention layer, which facilitates information exchange, and are immediately split back into their respective streams for all other operations. To further decouple their dynamics and directly support our training objective (described in Section 4.2), each stream receives its own distinct timestep embedding via adaptive layernorm (AdaLN) (Peebles & Xie, 2022). After traversing the shared MMDiT layers, the two streams are routed into their own modality-specific DiT blocks for several layers of fine-grained, specialized denoising. This final stage allows the vision pathway to focus on reconstructing a semantically consistent future embedding, while the action pathway refines the low-level motor control trajectories, thereby improving the joint modeling of both control and world dynamics.
@@ 4.2 Joint training algorithm
By sampling separate timesteps, we allow for causal dependencies to be learned between the modalities. For example, the model might be asked to predict a nearly-clean future observation ($\tau_{o}\approx 1$) from a completely noisy action ($\tau_{A}\approx 0$). This forces it to learn the inverse relationship, effectively answering: "What action must have been taken to achieve this future state?" Conversely, the model might be given a nearly-clean action ($\tau_{A}\approx 1$) and be required to denoise a very noisy future observation ($\tau_{o}\approx 0$). This trains the forward causal relationship: "What will the future state look like as a result of this action?" This varied training across all combinations of noise levels is what enables the model to capture the causal relationships between the modalities.
Decoupled noise scheduling.  Let the two modalities be the action chunk $A_{t}\in\mathbb{R}^{k\times d_{A}}$ and the future observation embedding $\tilde{o}_{t+k}\in\mathbb{R}^{d_{o}}$, where $d_{A}$ and $d_{o}$ are the dimensions of the action space and future image embedding, respectively. During training, we sample timesteps independently, with $\tau_{A}\in[0,1]$ for actions and $\tau_{o}\in[0,1]$ for future visual observations. Let $\epsilon_{A},\epsilon_{o}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ be sampled Gaussian noise, with which we noise $A_{t}$ and $\tilde{o}_{t+k}$, giving the noisy action sequences and noisy future observations as $A_{t}^{\tau_{A}}=\tau_{A}A_{t}+(1-\tau_{A})\epsilon_{A}$ and $\tilde{o}_{t+k}^{\tau_{o}}=\tau_{o}\tilde{o}_{t+k}+(1-\tau_{o})\epsilon_{o}$, respectively. The diffusion model $V_{\theta}$ predicts the velocity field of each modality, conditioned on the VLM feature $\Phi_{t}$. Let us denote $V_{\theta}(\Phi_{t},A_{t}^{\tau_{A}},\tilde{o}_{t+k}^{\tau_{o}},o_{t}^{s})=[V_{\theta}^{A},V_{\theta}^{o}]$ be the outputs of diffusion model. Then, the training objective for each action and image observations (i.e., world-modeling) are given as follows:
@@
where $\lambda_{\textrm{WM}}>0$ is a weighting hyperparameter for world-modeling loss.
Figure 3: Overview of vision-action joint sampling. During inference, we sample over $N_{A}$ steps for action tokens and $N_{o}=q\times N_{A}$ steps for vision tokens. The global timestep advances by $\Delta\tau_{o}=1/N_{o}$, where vision tokens are updated every step and action tokens are updated only every $q$ steps in $\Delta\tau_{A}=1/N_{A}$ strides. The default $q$ value is 1, and increasing it allows test-time scaling.